When Anders makes scones he uses sugar butter and flour in the ratio 1:2 8.
He has 100 gram of sugar and 300 gram of butter and 400 gram of flour in the kitchen cupboard.
Anders makes as many scones as he can with these ingredients.
How much of each ingredient does he use?

Answer by Theo(13342)

400 divided by 8 = flour for 50 scones.
300 / 2 = butter for 150 scones.
100 / 1 = sugar for 100 scones.
it looks like the constraint is flour.
the most number of scones he can make is 50 because of the flour limitation.
50 scones will require:
50 * 8 = 400 grams of flour.
50 * 2 = 100 grams of butter.
50 * 1 = 50 grams of sugar.